# Strong Coupling Constant Determination from the new CTEQ-TEA Global QCD Analysis

### v1

@@ V.2 Hierarchical Models via a Gaussian Mixture Model
where $w_{k}$ are mixture weights. The optimal value of $K$ is determined using model selection criteria such as the Akaike Information Criteria (AIC) [78] and Bayesian Information Criteria (BIC) [79]. For the present combination, the choice of $K=2$ minimizes both the AIC and BIC. Note that $K=1$ corresponds to the standard $\chi^{2}$ combination in Eq. (26). Minimizing the GMM log-likelihood yields a bimodal posterior for $\alpha_{s}(M_{Z})$. Following the procedure of Ref. [77], we extract the $\alpha_{s}(M_{Z})$ central value and its uncertainty from this distribution, and find
@@
We observe that, adopting a tolerance $T$ such that $\Delta\chi^{2}=T^{2}\simeq 33$ (equivalent to a scaling factor $e_{S}\simeq 6.2$, see Sec. VI.2 for the definition of the error scaling factor) produces an $\alpha_{s}(M_{Z})$ uncertainty from $\Delta\chi^{2}$ that is comparable to that obtained via the GMM:
$\alpha_{s}$ $=0.1183\pm 0.0023\quad(\text{scaled errors, }e_{S}\approx 6.24\text{ or }T^{2}\simeq 33).$
In contrast, a scaling value of $e_{S}\approx 6.9$ or equivalently, $T^{2}\approx 48.2$, produces the same value of uncertainty estimate as produced by the global tolerance criterion given in Eq. (29).
Figure 13 (right) illustrates the dependence of the $\alpha_{s}(M_{Z})$ uncertainty on the error scaling factor $e_{S}$ for both $K=1$ (standard $\chi^{2}$) and $K=2$ (GMM). The solid– and dashed–black horizontal lines indicate the values of $e_{S}$ defined above. The orange shaded region represents the uncertainty estimate using $K=2$, whereas the green region represents the uncertainty for $K=1$. When the errors are not rescaled, then $e_{S}=1$ and $K=1$ produces a very small uncertainty identical to the values presented in the last column of Table 4. On the other hand, the preferred model for $e_{S}=1$ corresponds to $K=2$. An increase in the value of $e_{S}$ corresponds to an increase in the uncertainty for $K=1$, whereas the uncertainty for $K=2$ decreases at $e_{S}\approx 3.5$. This is the region where the AIC and BIC criteria start to exhibit a preference for the $K=1$ model, cf. Sec V.3. It is interesting to note that one could in principle use this fact to determine a value of $e_{S}$ that reduces tension by an amount that makes the $K=1$ model preferable and in some sense making the data set more consistent. One could then use the uncertainty obtained in the $K=1$ case when $e_{S}\approx 3.5$, to determine the $\alpha_{s}(M_{Z})$ uncertainty. This choice however, produces poor fit with a large value of the reduced $\chi^{2}$ as well as a very small uncertainty, we therefore do not use this approach here.
V.3 Information Criteria and the GMM
Figure 14: Variation of the AIC with the error scaling factor $e_{S}$ shown for both $K=1$ (red) and $K=2$ (blue) in the GMM. Also shown are the values of $e_{S_{PDG}}\approx 4.7$ (dot-dashed) and $e_{S}\approx 6.2$ (dotted) which gives similar values of the uncertainty using the global tolerance criteria in Eq. 29.
We use the AIC to determine the optimal number of Gaussian mixtures $K$ for the GMM. The AIC is defined as
$\text{AIC}=N_{p}\ \ln N_{\textrm{pt}}-2\ln L(\bar{\alpha}_{s})$ (24)
where $L(\bar{\alpha}_{s})$ is the GMM likelihood given in Eq. 22 evaluated at the optimized value of $\alpha_{s}$, $N_{\textrm{pt}}$ is the number of data clusters and the number of independent parameters $N_{p}$ for the GMM is given by
$N_{p}=K\times N_{\text{parm}}+(K-1)\ .$ (25)
Since we profile the $\chi^{2}$ over $\alpha_{s}(M_{Z})$, $N_{\text{parm}}=1$. We vary the value of $K$ from $1$ through $5$ and find that the value of $K$ that minimizes the value of the AIC is $K=2$. We further explore the impact of inflating uncertainties through the error scaling factor ($e_{s}$) for two models: the usual $\chi^{2}$ in Eq. 26, which corresponds to the choice $K=1$ in the GMM, and $K=2$. The variation of AIC with $e_{s}$ for these two models are shown in Fig. 14. Without rescaling the errors, i.e for $e_{s}=1$, we find that the $K=1$ gives lower values of AIC and is therefore preferred. For this value of $e_{s}$, the $K=2$ model finds two possible values of $\alpha_{s}(M_{Z})=\left\{{0.1154,0.1193}\right\}$. As the value of $e_{s}$ increases the data becomes more consistent, thus increasing the value of AIC for $K=2$ and reducing the value of AIC for $K=1$. Beyond the value of $e_{s}\approx 2.3$, the $K=1$ model becomes the preferred choice. For $e_{s}\approx 2.3$, the fit is still of poor quality with a large reduced $\chi^{2}$, hence we do not use this as a measure to set the value of $e_{s}$. For values of $e_{s}\gtrsim 3.5$, we see that the curves for AIC for both models increase linearly at the same rate. For these value of $e_{S}$ the $K=2$ model does not find two numerically distinct values of $\alpha_{s}(M_{Z})$ and therefore begins to behave like the $K=1$ model, albeit some over-fitting due to the larger number of parameters. This also explains the kink like feature in Fig. 13 where there is a sudden reduction of the uncertainty for the $K=2$ model and the uncertainty for the $K=2$ model tracks the behavior of the $K=1$. The slightly larger uncertainties for $K=2$ for $e_{s}\gtrsim 3.5$ arise due to overfitting [77].
@@ VI Final combination of $\alpha_{s}$ determinations
In Sec. III.1 we pointed out that the $\Delta\chi^{2}=1$ criterion would lead to an unrealistically small $\alpha_{s}(M_{Z})$ uncertainty, and that an additional $\delta_{68\%}$ uncertainty must be associated to $\chi_{0}^{2}$. As a reminder, when examining the $\chi^{2}$ profile for the CT25prel baseline Fig. 4, the application of the $\Delta\chi^{2}=1$ criterion to the total $\chi^{2}$ (black curve) would result in $\alpha_{s}(M_{Z})=0.11837^{+0.00035}_{-0.00041}$. This uncertainty is much smaller than the global average quoted in the PDG [61].
@@
The limitation associated with a $\Delta\chi^{2}=1$ criterion and the significant discrepancies observed in the preferred $\alpha_{s}(M_{Z})$ value by different experiments, raise the question of finding consistent procedures to combine the $\alpha_{s}(M_{Z})$ values and their corresponding uncertainties. We address this question in this section.
@@ VI.1 The central $\alpha_{s}$ value
In Sec. II, we observed that the choice between the acceptable baseline data sets affects the central $\alpha_{s}$ value. The two described baselines, CT25prel and CT25, yield the central values of 0.1179 and 0.1185, respectively. In the following, we will quote the latter as our final $\alpha_{s}$ determination. In contrast, the choice between the baselines only weakly influences the uncertainty determination, which instead depends on the method for the estimation. In the following, we summarize the uncertainty ranges obtained with several such methods.
VI.2 Uncertainty with the Particle Data Group prescription
[…]
$\alpha_{s}(M_{Z})\times 10^{3}$ 115.14 118.62 120.65 118.37
Error $(\sigma_{i}\times 10^{3})$ 0.629 0.743 0.543 0.360
[…]
A combination of these determinations under the independence assumption requires the minimization of the following $\chi_{\alpha_{s}}^{2}$ log-likelihood function:
$\chi_{\alpha_{s}}^{2}=\sum_{i}\frac{(\alpha_{s}(M_{Z})-\alpha_{s,i}(M_{Z}))^{2}}{\sigma_{i}^{2}}\ ,$ (26)
where $\alpha_{s,i}$ and $\sigma_{i}$ denote the central value and uncertainty for the $i$-th dataset listed in Table 4. The combined estimate and its variance are given by
$\bar{\alpha}_{s}=\frac{\sum_{i}\alpha_{s,i}/\sigma_{i}^{2}}{\sum_{i}1/\sigma_{i}^{2}},\qquad\sigma^{2}=\frac{1}{\sum_{i}1/\sigma_{i}^{2}}\,$, (27)
which result in $\bar{\alpha}_{s}=0.1184\pm 0.0004$, whose mean is consistent with the value obtained from the global $\chi^{2}$ profile, thereby validating the underlying assumptions. However, the resulting uncertainty is significantly smaller than the uncertainties obtained in Sec. III.2 and than the current world average [61]. In addition, it produces $\chi^{2}/\text{d.o.f.}\approx 22$, indicating a dramatically bad quality of the fit. Therefore, this approach fails to capture the features of the various data in the baseline.
To improve on this, we apply the scale factor $e_{S_{\text{PDG}}}$ as recommended by the PDG [61], that increases the uncertainties in such a way that the reduced $\chi^{2}/\text{d.o.f.}\approx 1$. This condition is guaranteed by choosing the scale factor as
$e_{S_{\text{PDG}}}=\sqrt{\chi^{2}/\text{d.o.f.}}\ .$ (28)
$\alpha_{s}$ $=0.1184\pm 0.0017\quad(\text{scaled errors, }e_{S_{\text{PDG}}}\approx 4.7).$
It is important to point out that such a rescaling procedure is not advised and is unsuitable in fits with poor quality. In such cases, an alternative, informed estimate of the uncertainty should be used. Motivated by this, we investigate other different approaches to estimate the $\alpha_{s}(M_{Z})$ uncertainty.
VI.3 Final $\alpha_{s}$ determination
The four approaches that we choose for the final combination are based on the global and dynamical tolerances, as well as the two Bayesian models.
Global and dynamical tolerances. Sec. IV.1 described the commonly used uncertainty prescriptions based on the global and dynamical tolerance, or GT and DT. When we apply the GT criterion based on the 68% quantile for the sum of all $\chi^{2}_{E}$ from the individual experiments, we obtain
@@ VI.3 Final $\alpha_{s}$ determination
These results are obtained for the default pole mass $m_{c}=1.3$ GeV of the CT25 analysis. While in principle the charm mass is correlated with the gluon PDF and therefore $\alpha_{s}$, in practice this correlation is weak. Appendix B shows the GT and DT $\alpha_{s}$ ranges for the alternative masses 1.2 and 1.4 GeV, and those essentially coincide with the estimates for 1.3 GeV. The $m_{c}$ dependence is hence found to be minimal, as is also illustrated by the $\chi^{2}$ profiles in Fig. 15.
Bayesian Models. Section V deployed two Bayesian models to estimate $\alpha_{s}$ uncertainties. We obtained $\alpha_{s}(M_{Z})=\alpha_{s}=0.1181^{+0.0022}_{-0.0023}$ with the Bayesian Hierarchical Model and $\alpha_{s}(M_{Z})=0.1183\pm 0.0023$ with the Gaussian Mixture Model.
@@
Global Tolerance 29 ${0.1184}^{+0.0026}_{-0.0028}$
Dynamical Tolerance 30 ${0.1184}^{+0.0024}_{-0.0012}$
BHM 21 $0.1181^{+0.0022}_{-0.0023}$
GMM 23 $0.1183^{+0.0023}_{-0.0023}$
@@
Table 5 collects the uncertainty estimates obtained with our four methods. The last row, labeled “Average”, corresponds to approximating the $\alpha_{s}(M_{Z})$ probability distribution from each statistical model as a split normal distribution and summing over each model with uniform weights to produce an average. This outcome average value and its uncertainty correspond to $\alpha_{s}(M_{Z})={0.1183}^{+0.0023}_{-0.0020}$ .
Good consistency among the four determinations increases our confidence in the result. It is important to emphasize that the determinations do not invoke a fixed tolerance value, or $T^{2}$. Instead, they would justify a specific $T^{2}$, should it be used as a proxy for the more complex approaches that we describe. For the CT25 baseline, the average $\delta\alpha_{s}(M_{Z})$ in Table 5 translates into the $\Delta\chi^{2}={}^{+39}_{-25}$ for the global $\chi^{2}$. This is close, although not identical, to the approximate $T^{2}=37$ at 68% CL introduced in the “CT two-tier tolerance” prescription used in the previous CT PDF analyses [69, 80, 16].
@@ VII Conclusions
In this work, we presented a new determination of the strong coupling constant $\alpha_{s}(M_{Z})$ from the CT25 global QCD analysis at NNLO, incorporating high-precision LHC Run-2 data and improved methodologies. The final value of $\alpha_{s}(M_{Z})$ extracted from the CT25 global fit with its 68% CL uncertainty is
$\alpha_{s}(M_{Z})=0.1183^{+0.0023}_{-0.0020}.$ (31)
To assess in detail the variability of the central value and the $\alpha_{s}(M_{Z})$ uncertainty, we employed a suite of statistical models including global and dynamical tolerance criteria, Bayesian hierarchical models, and Gaussian mixture models. We pointed out that this determination is nominal, and we identified several assumptions that may modify the outcome.
The key challenge for increasing the accuracy of the $\alpha_{s}$ determination arises from insufficient information about the experimental correlated systematic errors that is made available to PDF fitters. In Sec. III.2, we investigated the sensitivity of $\alpha_{s}(M_{Z})$ to systematic uncertainty treatments in the data, particularly in inclusive-jet production, which shows significant variations in its preferences for the central value and uncertainty depending on the adapted treatment of its systematic errors. Two different choices for including the published systematic errors were considered: the multiplicative treatment normalizes the errors to the theory predictions; the additive treatment normalizes the error to the central value of the measurement. The multiplicative treatment produces results consistent with the world average, while applying the additive treatment across all jet measurements reveals substantial internal tensions among jet-production data sets. The actual nature of the correlated errors, which are likely to be neither fully multiplicative nor additive, may show deviations from these two scenarios. Experiments can reduce the associated uncontrolled ambiguity by publishing more complete models of their systematic uncertainties.
With the dynamic tolerance and cross-validation methods, the uncertainty estimates are sensitive to how the data is clustered. Section IV explored in detail the role of data clustering and random reshuffling of the residuals and nuisance parameters. We introduced a concept of data clustering safety that requires attention when tensions among data sets are present. The key idea is that the uncertainty on a fundamental parameter should not depend on how the individual data points are clustered, or at least the uncertainty associated with clustering must be estimated. Clustering safety is relevant both for the Hessian and Monte Carlo PDF methodologies and is more challenging when the experiments disagree.
In Sec. III.1, we also explored the impact of a large set of PDF parametrizations. In the case of the $\alpha_{s}$ determination, the impact of the parametrization is small, as demonstrated by the tight clustering of best-fit values across roughly 287 alternative models in Fig. 1. These parametrizations have the functional form $f(x,Q_{0})=x^{a}(1-x)^{b}P(x)$ where $P(x)$ is an $n$-th order polynomial in $x$ with $2\leq n\leq 5$.
The impact of charm-quark mass variations was found to be minimal, with negligible shifts in the central value and uncertainty of $\alpha_{s}(M_{Z})$.
Future experiments, seeking events coming from interactions beyond the standard model (BSM), will need to know trustworthy uncertainty ranges for the value of $\alpha_{s}(M_{Z})=\alpha_{s}(M_{Z})^{\textrm{(}central)}\pm\delta\alpha_{s}$.

### v2

@@ V.2 A Gaussian Mixture Model
where $w_{k}$ are mixture weights. The optimal value of $K$ is determined using model selection criteria such as the Akaike Information Criteria (AIC) [87] and Bayesian Information Criteria (BIC) [88]. For the present combination, the choice of $K=2$ minimizes both the AIC and BIC. Note that $K=1$ corresponds to the standard $\chi^{2}$ combination in Table 2. Minimizing the AIC score with respect to $K$ yields a bimodal posterior for $\alpha_{s}(M_{Z})$. Following the procedure of Ref. [86], we extract the $\alpha_{s}(M_{Z})$ central value and its uncertainty from this distribution and find
@@
Importantly, this estimate is obtained without inflating the errors, i.e., for the error rescaling factor $e_{S}=1$. We have also examined the combination under the tentative scaling of errors by a factor $e_{S}>1$. We find that the $K=2$ combination with $e_{S}\approx 1$ is always preferred (has the lowest AIC score), closely followed by the traditional $K=1$ combination with $e_{S}\approx 3.5$ that has a marginally higher score – see Fig. 14 in Appendix D. Thus, it would not be unreasonable to also consider the $K=1$ case when $e_{S}\approx 3.5$ to determine the $\alpha_{s}$ uncertainty. This choice, however, produces a poor agreement with the input data in Table 2, with a large value of $\sum_{i}\chi^{2}(\overline{\alpha}_{s},\alpha_{s,i},\sigma_{i})$ as well as a too small uncertainty. We therefore do not pursue it.
Figure 11 (right) illustrates the dependence of the $\alpha_{s}(M_{Z})$ uncertainty on $e_{S}$ for both $K=1$ (standard $\chi^{2}$) and $K=2$ (GMM). The orange and green shaded regions represent the uncertainty estimates for $K=2$ and $1$, respectively. When the errors are not rescaled, then $e_{S}=1$, and $K=1$ produces a very small uncertainty identical to the one presented in the last column of Table 2. On the other hand, for this $e_{S}$ value (specifically, for $e_{S}<2.3$), the $K=2$ model is preferred according to the AIC and BIC criteria. The uncertainty for $K=1$ increases with $e_{S}$, whereas the uncertainty for $K=2$ decreases around $e_{S}\approx 4.2$ and then increases linearily as in the $K=1$ model, see Figure 11.
From Fig. 11 (right), it follows that the standard ($K=1$) treatment reproduces the $\alpha_{s}$ uncertainty obtained with the GMM ($K=2$) by applying $e_{S}\approx 6.2$, or, equivalently, adopting a tolerance $T$ such that $\Delta\chi^{2}=T^{2}\approx 33$ (see Sec. IV.2 for the relation between $T$ and $e_{S}$):
$\alpha_{s}$ $=0.1183\pm 0.0023\quad(\text{scaled errors, }e_{S}\approx 6.2\text{ or }T^{2}\approx 33).$
This is larger than $e_{S}=4.7$ from the PDG prescription in Sec. IV.2 and slightly smaller than the effective scaling of $e_{S}\approx 6.9$, or $T^{2}\approx 48$, that reproduces the uncertainty estimate from the global tolerance criterion defined in Sec. IV.3. In Fig. 11 (right), the values of $e_{S}=4.7$ and 6.2 are indicated by vertical dot-dashed and dotted lines.
@@ VI Final combination of $\alpha_{s}$ determinations
In Sec. IV.1, we pointed out that the $\Delta\chi^{2}=1$ criterion would capture only a small part of realistic $\alpha_{s}$ uncertainty, and that an additional uncertainty must be associated with data’s internal consistency. As a reminder, when examining the $\chi^{2}$ profile for the CT25prel baseline Fig. 4, the application of the $\Delta\chi^{2}=1$ criterion to the total $\chi^{2}$ (black curve) would result in $\alpha_{s}(M_{Z})=0.11837^{+0.00035}_{-0.00041}$. This uncertainty is smaller than the global average quoted in the PDG [72].
@@
These limitations of the $\Delta\chi^{2}=1$ criterion raise the question of finding more complete procedures to quantify the uncertainty. We have explored several such procedures in Sec. IV and Sec. V. In this section we compile the results from these different prescriptions and combine them into one final estimate of the $\alpha_{s}$ central value and uncertainty.
@@ VI.1 The central $\alpha_{s}$ value
In Sec. II, we observed that the choice between the acceptable baseline data sets affects the central $\alpha_{s}$ value. The two described baselines, CT25prel and CT25, yield the central values of about 0.1179 and 0.1184, respectively. In the following, we will quote the latter as our final $\alpha_{s}$ determination. In contrast, the choice between the baselines only weakly influences the uncertainty determination, which instead depends on the method for the estimation. In the following, we summarize the uncertainty ranges obtained with several such methods.
VI.2 Final $\alpha_{s}$ determination
The four approaches that we choose for the final combination are based on the global and dynamical tolerances, as well as on the two Bayesian models.
Global and dynamical tolerances. Sec. IV.3 described the commonly used uncertainty prescriptions based on the global and dynamical tolerance, or GT and DT. When we apply the GT criterion based on the 68% quantile for the sum of all $\chi^{2}_{E}$ using the CT25 baseline from the individual experiments, we obtain
@@ VI.2 Final $\alpha_{s}$ determination
These results are obtained for the default pole mass $m_{c}=1.3$ GeV of the CT25 analysis. While in principle the charm mass is correlated with the gluon PDF and therefore $\alpha_{s}$, in practice this correlation is weak. Appendix E shows the GT and DT $\alpha_{s}$ ranges for the alternative masses 1.2 and 1.4 GeV, and those essentially coincide with the estimates for 1.3 GeV. The $m_{c}$ dependence is hence found to be minimal, as is also illustrated by the $\chi^{2}$ profiles in Fig. 15.
Bayesian Models. In Section V, we discussed two Bayesian models to estimate $\alpha_{s}$ uncertainties. We obtained $\alpha_{s}(M_{Z})=0.1181^{+0.0022}_{-0.0023}$ with the Bayesian Hierarchical Model and $\alpha_{s}(M_{Z})=0.1183\pm 0.0023$ with the Gaussian Mixture Model.
@@
Global Tolerance 28 ${0.1184}^{+0.0026}_{-0.0028}$
Dynamical Tolerance 29 ${0.1184}^{+0.0024}_{-0.0012}$
BHM 25 $0.1181^{+0.0022}_{-0.0023}$
GMM 27 $0.1183^{+0.0023}_{-0.0023}$
@@
Table 3 collects the uncertainty estimates obtained with our four methods. The last row, labeled “Average”, corresponds to approximating the $\alpha_{s}(M_{Z})$ probability distribution from each statistical model as a split normal distribution and summing over each model with uniform weights to produce an average. This yields a final combined average value and uncertainty of $\alpha_{s}(M_{Z})={0.1183}^{+0.0023}_{-0.0020}$ .
Good consistency among the four determinations increases our confidence in this result. It is important to emphasize that the determinations do not invoke a fixed tolerance value. Instead, they would justify a specific $T^{2}$, should it be used as a proxy for the more complex approaches that we describe. For the CT25 baseline, the average $\delta\alpha_{s}(M_{Z})$ in Table 3 translates into the $\Delta\chi^{2}={}^{+39}_{-25}$ for the global $\chi^{2}$. This is close, although not identical, to the approximate $T^{2}=37$ at 68% CL introduced in the “CT two-tier tolerance” prescription used in the previous CT PDF analyses [68, 89, 16].
@@ VII Conclusions
We have presented a new determination of the strong coupling constant $\alpha_{s}(M_{Z})$ from the CT25 global QCD analysis at NNLO, incorporating high-precision LHC Run-2 measurements and improved statistical methodology. The global analysis simultaneously fits parton distribution functions, $\alpha_{s}$, and other QCD parameters to a diverse collection of measurements from fixed-target and collider experiments, providing the most complete treatment of correlations among these parameters. Our central result, combining four independent uncertainty prescriptions, is
$\alpha_{s}(M_{Z})=0.1183^{+0.0023}_{-0.0020}\quad(68\%~\mathrm{CL}).$
This value is in good agreement with the current PDG world average and with other recent determinations from LHC data and lattice QCD.
A central theme of this work is the critical assessment of uncertainty quantification. We demonstrated that the naive $\Delta\chi^{2}=1$ uncertainty—approximately $\pm 0.0004$ for the CT25 data set—is insufficient, because it captures only statistical fluctuations and ignores the significant tensions among the three broad classes of data (DIS, Drell–Yan, and jet + $t\bar{t}$ production). These tensions are reflected in the large spread of preferred $\alpha_{s}$ values across process types (0.1152 to 0.1198 for the CT25prel baseline with multiplicative systematic errors) and in the failure of the combination of the three classes to satisfy a standard goodness-of-fit test by a wide margin.
We investigated four methodologically distinct approaches to incorporate the full uncertainty:
  • Global tolerance (GT): Based on the 68% quantile of the total $\chi^{2}$ distribution, yielding
    $\alpha_{s}(M_{Z})=0.1184^{+0.0026}_{-0.0028}.$
    It is trivially insensitive to data clustering.
  • Dynamical tolerance (DT): Constructed from quantiles of the individual experiment profiles, giving
    $\alpha_{s}(M_{Z})=0.1184^{+0.0024}_{-0.0012},$
    with a pronounced asymmetry. We showed that DT estimates depend on how experiments are grouped into clusters, and we introduced the concept of data-clustering safety to characterize this dependence: an uncertainty estimate on a fundamental parameter is clustering-safe if it remains stable under reasonable regroupings of the data. The GT satisfies this criterion by construction; the DT does not in general and requires supplementary tests. In the limit of a single cluster the DT reduces to the GT, while in the limit of many clusters it approaches the $\Delta\chi^{2}=1$ uncertainty.
  • Bayesian hierarchical model (BHM): Introducing a hyperprior that accounts for unmodeled inter-experimental spread, and averaging over the hyperprior parameter $0\leq\beta\leq 1$, yields
    $\alpha_{s}(M_{Z})=0.1181^{+0.0022}_{-0.0023}.$
  • Gaussian mixture model (GMM): Determining the data-adaptive hyperprior via information criteria, we find a bimodal posterior driven by multimodal structure in the data which yields
    $\alpha_{s}(M_{Z})=0.1183\pm 0.0023.$
The four methods are in good mutual agreement. Their uniformly weighted combination gives the final value quoted above, corresponding to an effective global tolerance of $\Delta\chi^{2}={}^{+39}_{-25}$, close to the $T^{2}=37$ used in previous CT PDF analyses.
We also identified systematic uncertainties that could shift the central value. The treatment of correlated systematic errors in inclusive-jet production is the most consequential: switching the normalization convention from multiplicative (ksys=1) to additive (k=2) for jet data alone lowers the global $\alpha_{s}$ by about 0.003, while simultaneously producing a large internal tension within the jet data sets and a substantially poorer goodness-of-fit. Since the major jet-energy-scale uncertainties at the LHC are regarded as multiplicative, the ksys=1 result is preferred; nonetheless, this exercise underscores the importance of clear documentation of systematic-error conventions by experimental collaborations. We also found that the choice of PDF parametrization and variations of the charm-quark pole mass between 1.2 and 1.4 GeV have only a small effect on $\alpha_{s}$, with the parametrization uncertainty confined to $\approx\pm 0.0005$ across 287 independent candidate parametrizations.
Looking ahead, future HL-LHC measurements will further constrain $\alpha_{s}$ and the PDFs, but the precision gains will be meaningful only if accompanied by robust uncertainty quantification. The concept of data-clustering safety introduced here provides a concrete, calculable diagnostic for the stability of uncertainty estimates based on tolerance or cross-validation methods, applicable to both Hessian and Monte Carlo PDF methodologies. Together with the Bayesian approaches discussed in this work, the proposed methodologies provide a framework for self‑consistent uncertainty quantification, which is essential for a precise determination of $\alpha_{s}$ from hadron‑collider data.